# Exploring Facial Biomarkers for Depression through Temporal Analysis of Action Units

### v1

@@ Abstract
Depression is characterized by persistent sadness and loss of interest, significantly impairing daily functioning and now a widespread mental disorder. Traditional diagnostic methods rely on subjective assessments, necessitating objective approaches for accurate diagnosis. Our study investigates the use of facial action units (AUs) and emotions as biomarkers for depression. We analyzed facial expressions from video data of participants classified with or without depression. Our methodology involved detailed feature extraction, mean intensity comparisons of key AUs, and the application of time series classification models. Furthermore, we employed Principal Component Analysis (PCA) and various clustering algorithms to explore the variability in emotional expression patterns. Results indicate significant differences in the intensities of AUs associated with sadness and happiness between the groups, highlighting the potential of facial analysis in depression assessment.
Index Terms:
depression, action units, emotion, facial analysis, biomarkers, clustering, time-series
@@ I Introduction
The World Health Organization (WHO) estimates that depression affects millions of people worldwide, making it one of the most common mental health conditions[2]. Depression is a condition that considerably reduces everyday operations and quality of life. It is characterized by continuous sorrow, lack of interest in activities, and a variety of mental and physical difficulties [3]. A timely and accurate diagnosis is essential for managing and treating conditions effectually. Traditional diagnostic techniques frequently depend on subjective self-reported questionnaires and clinical interviews like the Beck Depression Inventory (BDI) [11], the Hamilton Depression Rating Scale (HDRS) [4], and the Patient Health Questionnaire (PHQ-8 and PHQ-9) [5]. Several factors could influence these evaluations.
@@
Using facial expression analysis to find biomarkers linked to depression is one potential method. Facial expressions, which are governed by the Facial Action Coding System (FACS)’s definition of facial action units (AUs), are widely used markers of emotional states and can reveal important information about a person’s mental state [8]. The FACS, proposed by Ekman and Friesen [8], provides an exhaustive framework for categorizing facial movements associated with specific emotions. This system allows for the detailed analysis of facial action units, which are the fundamental actions of individual facial muscles [10]. Additionally, people with depression often exhibit a higher prevalence of sadness and lower frequencies of happiness compared to healthy individuals. However, there is a need for more refined and quantitative analyses to understand better the relationship between AUs, emotions, and depression.
@@ II Related Work
Jones et al. (2018) [6] conducted an extensive study on the facial expressions of depressed and healthy individuals using the FACS. Their analysis revealed that depressed patients exhibited shows higher frequencies of AU1 (inner brow raiser), AU4 (brow lowerer), and AU15 (lip corner depressor), which are commonly associated with sadness and distress. Also, the study found that these patients displayed lower frequencies of AU12 (lip corner puller), which is related to expressions of happiness.
@@
These studies collectively highlight the importance of specific facial AUs and emotions in identifying depression. They provide a foundation for our research, which further investigates the dominance of certain AUs and emotions in depression versus people without depression.
@@ III-D PCA Clustering
Various clustering methods, including K-means clustering, Agglomerative clustering, and Gaussian Mixture Model (GMM), was applied to these principal components. These methods enabled the identification of coherent clusters of facial expression patterns, facilitating an unsupervised understanding of emotional expression variability among participants.
@@ IV-A Descriptive Statistics
The bar plot in Figure 4 shows the distribution of participants in our study across three categories: Depressed, Healthy, and Sub-clinical. A smaller group of participants (9 individuals) were categorized as Sub-clinical, indicating that they exhibited some signs of depression but did not meet the full criteria for a depressed classification.
@@ IV-B Mean Intensity Comparison of the Dominant Action Units
The plots 5, 6 and 7 include a horizontal dashed line representing the overall mean intensity of happiness for both groups, as well as shaded areas depicting the standard deviation. These additions provide a clear visual representation of the variability and central tendency of the data.
@@
The analysis of mean intensity comparisons of key action units (AU1, AU4, AU15) in Table III reveals significant differences in facial expressivity between depressed and healthy patients. Depressed individuals consistently exhibit higher intensities of these action units, corresponding to expressions of sadness, distress, and concern. These findings support the hypothesis that specific facial expressions can serve as biomarkers for depression, providing a potential avenue for more objective and automated psychological assessments.
@@ IV-C Mean Intensity Comparison of Dominant Expressions: Sadness and Happiness
In this section, we analyze the mean intensity of dominant expressions, specifically sadness, and happiness, between depressed and healthy patients. This comparison highlights the differences in emotional expressivity between the two groups, providing further insight into their emotional states.
@@ IV-D2 Clustering Analysis
In Figure 13 & 14 we see the results of Agglomerative clustering applied to the transformed data. The clusters are well-separated from each other, data points in one cluster are distinctly different from those in neighboring clusters where the cluster for healthy patients indeed contains significantly fewer points than the other.

### v2

@@ Abstract
Depression is characterized by persistent sadness and loss of interest, significantly impairing daily functioning and, now a widespread mental disorder. Traditional diagnostic methods rely on subjective assessments, necessitating objective approaches for accurate diagnosis. Our study investigates the use of facial action units (AUs) and emotions as biomarkers for depression. We analyzed facial expressions from video data of participants classified with or without depression. Our methodology involved detailed feature extraction, mean intensity comparisons of key AUs, and the application of time series classification models. Furthermore, we employed Principal Component Analysis (PCA) and various clustering algorithms to explore the variability in emotional expression patterns. Results indicate significant differences in the intensities of AUs associated with sadness and happiness between the groups, highlighting the potential of facial analysis in depression assessment.
@@ I Introduction
The World Health Organization (WHO) estimates that depression affects millions of people worldwide, making it one of the most common mental health disorders. Depression is a condition that considerably reduces everyday operations and quality of life. It is characterized by continuous sorrow, lack of interest in activities, and a variety of mental and physical difficulties [1]. A timely and accurate diagnosis is essential for managing and treating conditions effectually. Traditional diagnostic techniques frequently depend on subjective self-reported questionnaires and clinical interviews like the Beck Depression Inventory (BDI) [2], the Hamilton Depression Rating Scale (HDRS) [3], and the Patient Health Questionnaire (PHQ-8 and PHQ-9) [4]. Several factors could influence these evaluations.
@@
Using facial expression analysis to find biomarkers linked to depression is one potential method. Facial expressions, which are governed by the Facial Action Coding System (FACS)’s definition of facial action units (AUs), are widely used markers of emotional states and can reveal important information about a person’s mental state [7]. The FACS, proposed by Ekman and Friesen [7], provides an exhaustive framework for categorizing facial movements associated with specific emotions. This system allows for the detailed analysis of facial action units, which are the fundamental actions of individual facial muscles [5]. Additionally, people with depression often exhibit a higher prevalence of sadness and lower frequencies of happiness compared to healthy individuals. However, there is a need for more refined and quantitative analyses to better understand the relationship between AUs, emotions, and depression.
@@ II Related Work
Jones et al. (2018) [8] conducted an extensive study on the facial expressions of depressed and healthy individuals using the FACS. Their analysis revealed that depressed patients exhibited higher frequencies of AU1 (inner brow raiser), AU4 (brow lowerer), and AU15 (lip corner depressor), which are commonly associated with sadness and distress. Also, the study found that these patients displayed lower frequencies of AU12 (lip corner puller), which is related to expressions of happiness.
@@
These studies collectively highlight the importance of specific facial AUs and emotions in identifying depression. They provide a foundation for our research, which further investigates the dominance of certain AUs and emotions in depression versus participants without depression.
@@ III-D PCA Clustering
Various clustering methods, including K-means clustering, Agglomerative clustering, and Gaussian Mixture Model (GMM), is applied to these principal components. These methods enabled the identification of coherent clusters of facial expression patterns, facilitating an unsupervised understanding of emotional expression variability among participants.
@@ IV-A Descriptive Statistics
The bar plot in Figure 4 shows the distribution of participants in our study across three categories: Depressed, Healthy, and Sub-clinical. A smaller group of participants (9 individuals) is categorized as Sub-clinical, indicating that they exhibited some signs of depression but did not meet the full criteria for a depressed classification.
@@ IV-B Mean Intensity Comparison of the Dominant Action Units
The plots 5, 6 & 7 include a horizontal dashed line representing the overall mean intensity of happiness for both groups, as well as shaded areas depicting the standard deviation. These additions provide a clear visual representation of the variability and central tendency of the data.
@@
The analysis of mean intensity comparisons of key action units (AU1, AU4, AU15) in Table III reveals substantial differences in facial expressivity between depressed and healthy patients. Depressed individuals consistently exhibit higher intensities of these action units, corresponding to expressions of sadness, distress, and concern. These findings support the hypothesis that specific facial expressions can serve as biomarkers for depression, providing a potential avenue for more objective and automated psychological assessments.
@@ IV-C Mean Intensity Comparison of Dominant Expressions: Sadness and Happiness
In this section, we analyze the mean intensity of dominant expressions, specifically sadness and happiness, between depressed and healthy patients. This comparison highlights the differences in emotional expressivity between the two groups, providing further insight into their emotional states.
@@ IV-D2 Clustering Analysis
In Figure 13 & 14, we see the results of Agglomerative clustering applied to the transformed data. The clusters are well-separated from each other, data points in one cluster are distinctly different from those in neighboring cluster where the cluster for healthy patients indeed contains significantly fewer points than the other.
@@ V Conclusion
Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—SFB 1483—Project-ID 442419336, EmpkinS.